OpenStudy (anonymous):

find the locus of the point P(x,y) which moves so that its perpendicular distance from 3x+4y=0 is always equal to its perpendicular distance from 5x-12y=0

OpenStudy (anonymous):

Do i find the perpendicular distance of both sides, and set the two to equal? and then solve from there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{|3x+4y|}{\sqrt{3^2+4^2}}\]=\[\frac{|5x-12y|}{\sqrt{5^2+(-12)}}\]

OpenStudy (anonymous):

Great!....Thanks again

OpenStudy (anonymous):

(-12)^2

OpenStudy (anonymous):

you're welcome!